### v2

Abstract
Genomic studies, including CRISPR-based PerturbSeq analyses, face a vast hypothesis space, while gene perturbations remain costly and time-consuming. Gene expression models based on graph neural networks are trained to predict the outcomes of gene perturbations to facilitate such experiments. Active learning methods are often employed to train these models due to the cost of the genomic experiments required to build the training set. However, poor model initialization in active learning can result in suboptimal early selections, wasting time and valuable resources. While typical active learning mitigates this issue over many iterations, the limited number of experimental cycles in genomic studies exacerbates the risk. To this end, we propose graph-based one-shot data selection methods for training gene expression models. Unlike active learning, one-shot data selection predefines the gene perturbations before training, hence removing the initialization bias. The data selection is motivated by theoretical studies of graph neural network generalization. The criteria are defined over the input graph and are optimized with submodular maximization. We compare them empirically to baselines and active learning methods that are state-of-the-art on this problem. The results demonstrate that graph-based one-shot data selection achieves comparable accuracy while alleviating the aforementioned risks.
Keywords:
Active Learning Graph Neural Networks Gene Perturbation Prediction
1 Introduction
Genomic research enables the study of genetic factors underlying various health conditions, opening new avenues for therapeutic development. Techniques such as CRISPR interference (CRISPRi) and PerturbSeq [5] have revolutionized the landscape of genomic experimentation by enabling high-throughput screenings. However, analyzing combinatorial gene perturbations remains challenging [38, 28], leading most studies to focus on single-gene perturbations. Moreover, most CRISPR-based PerturbSeq experiments are restricted to the order of a few hundreds of gene perturbations [10, 1, 31] due to budget and time constraints, despite the availability of over 20,000 potential gene targets. To complement experimental observations, several machine learning models have been developed to predict the outcome of gene perturbations on a given cell  [34, 7, 8, 13, 22, 23, 14].
(a) Active learning.
(b) One-shot data selection.
Figure 1: The difference between active learning and one-shot data-selection for genomic experiments with GEARS [34].
These models leverage diverse neural architectures, ranging from autoencoders [7] to graph neural networks [34], and they require a large number of training samples to generalize effectively. This reliance on supervision occasionally defeats their main purpose (i.e., minimizing the experimental cost), because the labels are themselves retrieved from genomic experiments. Additionally, disease-specific variations in gene expression make it difficult to train a universal model, necessitating the development of models tailored to specific experimental conditions.
Active learning [37] offers a promising solution by enabling iterative interaction between the wet lab and the machine learning models, to optimize data collection and hence training, with a limited number of cycles, as shown in Fig. 1(a). In the context of CRISPRi-based PerturbSeq experiments, these cycles tend to be restricted (e.g. 5 cycles), as each experiment in the lab may take up to a month [17]. Given the vast number of possible genes to explore, active learning strategies are well-suited to efficiently train predictive models while minimizing the experimental costs.
However, active learning depends on the model’s initialization, which poses a non-negligible risk in this context. Fig. 2 illustrates the amount of variation induced in the genomic perturbations selected to train the GEARS model [34], based on 5 different random initializations. In particular, it illustrates the number of genes that are selected in any one given cycle as black bars and the total number of unique genes that are selected for the 5 differently initialized models as the orange bars. Since the orange bars are often more than twice as large than the black ones, we observe that the perturbed genes differ significantly for each initialization, i.e., an extreme sensitivity of the selected genes to the model’s initialization. The effect is more pronounced in the initial cycles, as the models become more stable in subsequent cycles and converge to similar choices. However, given the aforementioned limitation in the cycle horizon, the train set may mistakenly include gene perturbations that are difficult to correct afterwards. In typical active learning scenarios, a longer horizon leads to an overall stable selection, mitigating early-stage errors over time. Nevertheless, in this case, the wet lab may risk losing months of work and resources for a model that is not properly initialized and may not converge in time.
Figure 2: The total number of unique genes perturbed in active learning experiments in two datasets with 5 different random seed initializations of GEARS [34], on the same train-test split. Compared to the budget for gene perturbations, i.e., the number of genes to be perturbed in each cycle, this indicates the variation associated with building the dataset based on the model in the initial cycles.
To mitigate this risk, we explore the use of one-shot data selection methods as an alternative to active learning for the problem of training gene expression predictors. By selecting the samples upfront based on the input of the model, as seen in Fig. 1(b), we can provide competitive results while exhibiting:
  • Independence of model initialization. A data-selection method runs before the model’s training and hence it does not depend on the model’s predictions. This reduces the risk of wasting experimental resources on erroneous gene perturbations.
  • Reduced operational delay. Each iteration in active learning introduces extra computational overhead: model retraining, recalculating uncertainty scores, etc. Moreover iterative experiments exhibit delays due to logistical complexities and communication, compared to one large experiment. These problems can add a substantial delay, which is not present in one uniform experiment.
We propose two one-shot data selection methods for training the state-of-the-art GEARS model [34], although our methods can work with any graph-based predictor. We rely on theorems for GNN generalization to define proxy criteria to guide the sample selection based on GEARS’ input knowledge graph. The criteria are optimized greedily to build the train set. Our contributions are as follows:
  • Two novel data-selection criteria based on generalization theorems for GNNs that utilize the input knowledge graph and are model-free.
  • A cost effective lazy forward submodular maximization algorithm to optimize one of the criteria, and a respective greedy heuristic for the other.
  • An empirical comparison including two active learning methods (one of which is SOTA in our problem) and two baselines, indicating the effectiveness of one-shot data selection.
2 Related Work
The body of related work can be separated in two categories. The first one aims to find the optimal gene perturbations of an experiment such that a causal quantity is maximized (e.g., a phenotype). The second studies efficient training strategies for models that provide gene expression predictions for any given cell. We will try to delineate the two research veins to clarify the position of our approach in the literature.
The problem of optimal experimental design consists of a large amount of treatment options (i.e., gene perturbations) and the causal outcome variable (e.g. increase in guide RNAs or T cells [26]) measured on average over the experimental samples. The task is to find a way to maximize the causal outcome using the treatments and a fixed number of interventions (experiment cycles) with experimental outcomes, i.e., accelerating the discovery of useful perturbations that change a given phenotype. This can be applied to CRISPRi assays where the cheap and fast nature of the interventions allows for a large number of experimental cycles, while the quantity maximized is the causal outcome (typically a scalar) related to the cell. In static optimal experimental design, the treatment allocation is specified beforehand while in the sequential setting, the design adapts to the output of a prior experiment [11]. Bayesian Optimization or online learning are popular solutions to this problem  [30, 24]. For example, in DiscoBax [24] or OAE [30], the reward is defined by the expected value of the phenotype and the evaluation is based on how many of the best interventions are recalled [26].
In contrast to optimal experimental design, active learning aims to train effectively a model that can predict the gene expression after a cell has undergone a specific gene perturbation. The base models, such as GEARS [34], can be used to predict directly the outcome for certain perturbations. The function we aim to maximize is not an experimental metric based on a hypothesis, but rather the accuracy of the model’s prediction in an unseen test set. This is useful in PerturbSeq experiments, which combine CRISPR and RNA sequencing, where experimental cycles are more costly and time-consuming and the gene expressions are the experimental outcomes. Computationally this is a multidimensional regression problem where we can choose which new samples to label to improve the model.
The closest method to our approach is IterPert [17], which uses active learning and prior multimodal knowledge to build a train set for GEARS. The number of available rounds is diminished (5 rounds) compared to works like GeneDisco [26], since each round can take close to a month in the wet-lab in a PerturbSeq experiment, rendering a bigger horizon infeasible. The problem is addressed from the perspective of active learning under budget. Instead of targeting diverse samples expected to increase the model’s generalization, the rudimentary sample selection is preferred [15]. The method outperforms numerous active learning techniques, albeit with the use of prior knowledge. In fact, prior knowledge incorporation is so effective, that it suffices to outperform the active learning methods without any updates, i.e., the model IterPert-prior-only. However, prior knowledge is not guaranteed to exist for every dataset.
Our one-shot data-selection methodology addresses this gap along with the fundamental problems of the active learning analyzed in sec. 1. We focus on training GEARS effectively while lacking prior information and the ability to take feedback from the model. The former is a plausible assumption since data for specific hypothesis are not that commonly available, and the latter is motivated by the variation inherent in the initial cycles of an active learning process and the cost to acquire samples in our context.
3 Methodology
In this section we first formulate the problem and clarify the motivation for pure graph-based methods for one-shot data selection. Subsequently we introduce the two methods for one-shot data selection.
3.1 Formulation
We are given a gene graph $G=(V,E),$ with $|V|=N$ and $|E|=M$, where some of the nodes are candidate perturbations. We assume that the dataset is partially labeled, meaning there is a subset of nodes $\mathcal{L}\subset V$ labeled. Each label is a matrix $Y_{u}\in\mathcal{R}^{c\times|g|}$ which represents the gene expression of a set of genes $g$ (which can belong to $V$ or not) of $c$ cells that underwent perturbation $u\in V$, along with the respective pre-perturbation gene expressions $X_{u}\in\mathcal{R}^{c\times g}$. There is respectively a set of genes $\𝑈\𝑉U=V\backslash\mathcal{L}$ that are unlabeled. GEARS learns how to predict $Y_{u}$ based on a given perturbation $u$. The problem we focus on is:
    Iteratively build the train set $S$ with the $b$ best genes to perturb from set $U$ to maximize GEARS’ accuracy on an unseen set of gene perturbations.
We aim to build the train set $S$ in order to cover as much and as diverse information as possible to train the model. Similar to the core-sets [36], we define a covering function that maximizes the model’s reach, but instead of defining it in the embedding space, we define it in the knowledge graph. The main reason behind this is that our graphs are not accompanied by node features. The node embeddings are randomly initialized and learned through training. This means any density-based method would be conditioned on the model initialization problem we explained in sec. 1.
To be specific, in GEARS and most gene expression models, the pre-perturbation gene expressions $X_{u}$ are not utilized during the prediction process as node features, but rather they are added to the predicted perturbation to arrive at $Y_{u}$. This happens because the prediction of the perturbation should be independent to the state of the cell $X_{u}$, as in real experiments, the cells will not exist until we have already decided which perturbation to run. This is why gene expressions models predict directly the gene expression displacement given the name of the perturbation, without using the cell $X_{u}$ [34, 7, 13]. Hence, in the absence of input node features and with the intention to remove the dependence of the train set selection on the model (due to the aforementioned variability in the resulting election), we are left with the knowledge graph as our sole remaining source of information to select the train set. The GEARS architecture is described in detail in the appendix for reference and further clarification.
3.2 GraphReach
To understand how to use the graph to guide our selection, we start with the graph convolution layer of GEARS. The Simplified Graph Convolution (SGC) layer [43] is defined as
$\tilde{\mathbf{A}}=\mathbf{D}^{-1/2}\mathbf{A}\mathbf{D}^{-1/2}$ (1)
$\mathbf{H}_{t+1}=\tilde{\mathbf{A}}^{K}\mathbf{H}_{t}\mathbf{W}$ (2)
where $\mathbf{H}_{t}$ are the node representations at layer $t$, $\mathbf{W}$ represents the learnable parameters of the model, $\mathbf{A}$ is the graph’s adjacency matrix, $K$ is the power of the adjacency predefined as 1 in GEARS, and $\mathbf{D}$ is the diagonal degree matrix. It should be noted that SGC is known to exhibit better generalization gaps than most of the other layers [18, 41], rendering it a well-suited module for our task.
Intuitively, our main goal is to label nodes such that we cover as much information as possible from the input, akin to core-sets. As mentioned above, the SGC aggregates information from the whole network; yet only the representations of nodes that are within the reach of labeled training nodes will get updated through the message passing layers. For example, if there are $L$ layers, message passing reaches up to the $L$-hop neighborhood of labeled nodes in set $S$. These are the only nodes relevant to training, i.e., their embeddings are updated based on supervision signal, as the rest of the graph does not effectively contribute to the training process. It is important to highlight this semi-supervised setting allows for the update of test nodes’ representations if they are sufficiently close to the training nodes. This clarifies how covering bigger parts of the graph can potentially improve the prediction.
Thus, to maximize the reach of the supervision signal and cover as much of the input information as possible, we should aim to maximize the number of nodes reached by the training set $S$. Beyond our intuition, this is supported by theoretical insights into the generalization capability of GNNs. The number of nodes that we have to predict and the number of samples that we can add in each cycle is fixed, so we can not change the VC dimension [35]. However, theoretical studies suggest that the generalization of a GNN is improved based on the distance between the training and the test nodes in the graph [25, 46], supporting our intuition that covering a larger part of the graph with training nodes-and thereby reducing the distance to the test set- is beneficial to the model. Note that these works define the distance as the structural and distributional difference; nevertheless, since we lack node features, the structural distance is the sole factor in our case.
Based on the above, we can define our criterion function, similar to an acquisition function in active learning, such that we maximize the number of nodes reached by the set $S$. Specifically for a new node $v$, the criterion is defined as the additional reachability that $v$ adds in the set of train perturbations $S$
$\alpha(v,S)=|R_{L}(S\cup\{v\})-R_{L}(S)|$ (3)
where $R_{L}$ is the set of nodes reachable from set $S$, based on the number of SGC layers $L$ in the definition of GEARS
$R_{L}(S,A)=\{v\in V\mid(\mathbf{A}^{\ell}_{uv}>0\text{ for some }u\in S,\text{and }1\leq\ell\leq L\}$ (4)
The subscript $L$ is omitted in subsequent discussion as it remains stable throughout all experiments in GEARS. The criterion function selects greedily the node $v$ maximizing $\alpha(v,S)$ in every iteration. This is fundamentally suboptimal [6, 20] but many active learning methods [24, 30, 42, 40] prove the base function is monotonic and submodular to achieve a guarantee that the result will be at least $1-1/e$ close to the optimal [29]. We follow suit and prove the same for reachability.
Lemma 1
The reachability function $R$ as defined in eq. 4 is monotonic and submodular.
Let two sets $A\subseteq B\subseteq V$ and any node $v\notin B$. Since $A\subseteq B$, every node reachable from $A$ is also reachable from $B$, meaning $R(A)\subseteq R(B)\rightarrow|R(A)|\leq|R(B)|$ which proves monotonocity. By definition, adding a node earlier ($\alpha(v,A)$) adds at least as many reachable nodes as adding it later ($\alpha(v,B)$), since the nodes covered between $A$ and $B$ can belong to $R(v)$ as well: $R(v)\cap(R(B)\setminus R(A))\geq 0$. This means that any node covered in $(R(B)\setminus R(A))$ is removed from $R(v)$, hence diminishing the additional reachability that $v$ can bring to $B$ compared to $A$. From this sketch we can contend that the reachability is submodular. The reader is referred to the appendix for the proof.
The greedy approach requires testing every available perturbation for every addition during a cycle, i.e., $\mathcal{O}(|U|T)$ where $T$ is the total training budget in samples. We employ a cost effective lazy forward approach to accelerate the acquisition step without loss of accuracy [21]. The final data selection algorithm, called GraphReach from graph reachability, is shown in Alg. 1.
Algorithm 1 GraphReach
1:budget $B$, graph $G$
2:Train set $S=\emptyset$
3:Criterion $\Delta[v]=\alpha(v,\emptyset)$ for all $v\in V$
4:sort($\Delta$)
5:Heap $\mathcal{L}$ with gene perturbations prioritized by $\Delta$
6:while $\mathcal{L}\neq\emptyset$ and $|S|\leq B$ do
7:     $v$ = $\mathcal{L}[0]$
8:     $u$ = $\mathcal{L}[1]$
9:     $\delta=\alpha(v,S)$
10:     if $\delta>\Delta[u]$ then
11:         $S=S\cup\{v\}$
12:         $\mathcal{L}=\mathcal{L}[1:]$
13:     else
14:         $\Delta[v]=\delta$
15:         sort($\Delta$) and rearrange $\mathcal{L}$ accordingly
16:     end if
17:end while
18:$S$
3.3 MaxSpec
Another line of research suggests that GNN generalization improves if the largest singular value (i.e., spectral norm) of the message passing operator is diminishing [18]. In our context, we can translate this as choosing a perturbation such that the underlying graph formed by the nodes that are reachable from $S$, has the smallest possible spectral norm of the respective diffusion matrix. The diffusion matrix is $\mathbf{D}^{-1/2}\hat{\mathbf{A}}_{S}\mathbf{D}^{-1/2}$ with $\hat{\mathbf{A}}_{S}=\mathbf{A}_{S}+\mathbf{I}d$ and $\mathbf{A}_{S}=A[R(S),R(S)]$ is the subgraph induced by $R(S)$. Hence we define a new criterion function similar to eq. 3, but instead of reachability, we utilize the spectral norm as:
$M(\mathbf{A}_{S})=\|\mathbf{D}^{-1/2}\hat{\mathbf{A}}_{S}\mathbf{D}^{-1/2}\|_{2}$ (5)
We keep using a greedy selection as a heuristic. We still rely on a cost-effective lazy forward method, hence the algorithm MaxSpec (from maximizing spectral norm) is the same alg. 1, but substituting the reachability function $R$ with $M$.
4 Experiments
In this section we first describe the design of our experiments, including the graph, the gene perturbation datasets, the evaluation methods, and the benchmarks used for comparison. Afterwards, we showcase the performance of the methods and analyze them to derive conclusions about the soundness of the proposed techniques.
4.1 Design
We test our methods in two genomic datasets stemming from PerturbSeq experiments, single-cell CRISPR-based screenings on K562 chronic myeloid leukemia cells:
  • Replogle: A study that systematically maps genotype-phenotype relationships using CRISPR knockouts. After data curation, it includes 1087 binary gene perturbations with approximately 150 cells each [33].
  • Adamson: A study that investigates the mammalian unfolded protein response of K562 cells to CRISPR interference such as repression. After data curation, it contains 81 binary gene perturbations with approximately 800 cells each [1].
The datasets are diverse in terms of supervision (150 to 800 cells) and number of perturbations (1087 to 82). We chose them in order to juxtapose the method’s accuracy in these two different settings and identify which methods perform adequately in both, taking into consideration the scarcity of relevant datasets.
We now describe the graph underlying both datasets. The graph we focus on, akin to [34], is based on pathway information from GO51 [9]. Each gene is associated with a number of pathway GO terms. The Jaccard similarity between the sets of pathways of two genes, i.e., the fraction of shared pathways, is used to calculate the strength of the edge between them. The graph is sparsified by keeping a predefined number of the most important neighbors for each node. The final graph contains over 9,800 nodes and 200,000 edges.
It should be noted that GEARS utilizes the whole graph to learn representations and uses the gene perturbation datasets for supervision only. In other words, the model uses the perturbations corresponding to the train set to compute the loss, but the SGC layer runs over the whole graph, including the test perturbations and unrelated genes, in a semi-supervised learning fashion [19]. The data selection methods follow suit and utilize the whole graph. For example the reachability of a gene perturbation in Replogle is not defined only on the 1081 perturbations, but on the 9,800 nodes.
Similar to  [17], we start with a random test set that contains roughly 10% of the available perturbations, i.e., 100 for Replogle and 8 for Adamson. The test set stays constant throughout the cycles to ensure there is no interaction between the strategies and the evaluation, which could lead to inconsistencies, e.g., adding a test perturbation that is hard to predict in the train set would diminish the error due to easier evaluation not due to the strategy’s success. We train the model 5 times sequentially, each time adding to the test set a fixed amount of gene perturbations (100 for Replogle and 8 for Adamson). This means that at the end of the process, around 40% of the dataset will not be utilized in training nor testing. This serves to make a stricter comparison of the selected train sets, meaning if we utilize the whole dataset by the end the methods would be indistinguishable.
We repeat the experiments 10 times for different random initializations (including models and data splits) and report the resulting standard deviation. The GEARS parameters are the defaults from the original code [^1] https://github.com/snap-stanford/GEARS/tree/master. The evaluation metrics (MSE and Pearson correlation) are computed on the top 20 most differentially expressed genes, i.e., the genes that exhibit the biggest expression differences in the experiment, akin to the literature [34, 17]. This is a common practice since the vast majority of the genes have zero expression, meaning the results would be skewed and the differences minuscule had we computed them globally. The code to reproduce the experiments is in an anonymized repository in the appendix.
The benchmark models are:
  • Random chooses randomly the set from the available gene perturbations, as in normal training.
  • Uncertainty utilizes GEARS’ inherent uncertainty estimation to choose the perturbations that exhibit the biggest uncertainty in every acquisition step.
  • ACS-FW [32] is a Bayesian batch active learning model. It selects perturbations such that the new posterior distribution of the model’s parameters approximates the expected posterior when the whole dataset is available. The problem is formulated as a sparse subset approximation and is solved with Frank-Wolfe optimization. We utilized the version from the BMDALreg library [16], which was one of the top performing methods in similar experiments on active learning for GEARS [17].
  • TypiClust [15] is the state of the art active learning algorithm on our problem, as it has outperformed a plethora (8 to be precise) of models, including BADGE [3], core-set [36] and ACS-FW in the aforementioned experiments on active learning for GEARS [17]. The algorithm clusters the candidate perturbations based on the final graph layer representation from GEARS. Within each cluster, the typicality is quantified as the inverse of the average distance between each sample and an example’s K-nearest neighbors, with K=20. The most typical sample is selected. Note again that the aforementioned IterPert does not operate without the use of prior-knowledge and hence can not be utilized in our setting. We utilized TypiClust, which is the next best method that relies only on the graph and the model.
4.2 Results
As can be seen in Tab. 1, the models developed from the data selection methods perform better than SOTA in one dataset and close to it in another. To be precise, while MaxSpec and TypiClust are the top methods in Reglogle and Adamson respectively, GraphReach has overall the most balanced performance being first or second in all datasets and metrics with minimal differences (considering the standard deviations) from the top.
Fig. 3(a) and 4(a) show that GraphReach outperforms marginally TypiClust in 3 out of 5 cycles in the Adamson dataset. Both achieve a better accuracy by the third cycle than the other models do in 5 cycles, with the exception of ACS-FW at the 5th cycle. The large standard deviations in Adamson stem from the reduced number of perturbations, which causes different test splits to exhibit significant variation in difficulty level. These variations are not directly associated with the variations from model initialization shown in Fig. 2, as they stem from the difference in data splits.
Category Strategy MSE ($10^{3}$) ↓ Pearson ($10^{3}$) ↑
Replogle Adamson Replogle Adamson
Random $156\pm 5$ $154\pm 44$ $919\pm 8$ $959\pm 9$
Active Learning Uncertainty $169\pm 5$ $170\pm 44$ $917\pm 8$ $954\pm 8$
ACS-FW $215\pm 72$ $146\pm 40$ $899\pm 20$ $960\pm 9$
TypiClust $155\pm 6$ $\mathbf{138\pm 47}$ $919\pm 7$ $\mathbf{963\pm 9}$
Data Selection MaxSpec $\mathbf{150\pm 4}$ $160\pm 49$ $920\pm 9$ $958\pm 11$
GraphReach $151\pm 4$ $143\pm 44$ $\mathbf{921\pm 8}$ $962\pm 9$
Table 1: Test mean squared error and Pearson coefficient of the top 20 most differentially expressed genes, over 10 random seeds (data split and model initialization) with standard deviations, at the end of a 5-cycle experiment with GEARS. Best method denoted with bold and second best is underlined.
For the Replogle dataset, the active learning methods are outperformed clearly by the graph-based data selection methods and even by random selection in the initial cycles. ACS-FW in particular has considerably worse performance than random, which is why it is not included in Fig. 3(b) and Fig. 4(b), while TypiClust consistently falls behind the remaining methods. In contrast, MaxSpec is the best method in Replogle but argueably the worst in Adamson. We hypothesize that this performance difference is related the drastic difference in size of these two datasets with Replogle containing several hundreds of nodes, whereas Adamson only considers graphs composed of tens of nodes. GraphReach on the other hand, is the only consistent method in both datasets from all benchmarks. This happens because optimizing for reachability has the same effect independently of the graph’s scale: maximizing the coverage of the train set based on the input information.
(a) MSE for Replogle.
(b) MSE for Adamson.
Figure 3: MSE error through the acquisition cycles (avg on 10 random seeds).
(a) Pearson for Replogle.
(b) Pearson for Adamson.
Figure 4: Pearson correlation through the acquisition cycles (avg on 10 random seeds).
Regarding the course of the models’ accuracy throughout cycles, it is visible that adding wrong perturbations actually deteriorates the accuracy of the model instead of keeping it stable e.g. cycle 3 of TypiClust in Replogle or cycle 4 of ACS-FW in Adamson. This indicates that the base model can overfit significantly in case of wrong additions. In a realistic setting with limited gene perturbations for training and a much larger test set (e.g. all available genes), we expect this effect to be greater. This further supports the need for the data selection methods, which have no cycle of deteriorating accuracy.
Additionally, the results are aligned with the literature that confirms random selection is stronger than most active learning in limited cycles [39, 4, 27, 47]. We observe that simple randomness outperforms the rest of the methods in 1 cycle in Adamson and Replogle, while being close in the first three iterations. However, in the final cycle, which is the main outcome of the study, random selection is clearly surpassed by more sophisticated methods.
Furthermore, we observe the ineffectiveness of Uncertainty as a standalone active learning strategy. On the one hand, this stems from the well-known open problem of quantifying uncertainty in deep learning [12, 2]. On the other hand, it is another facet of the disadvantages of active learning in limited cycles. Due to the model’s lack of training and the random initialization, the uncertainty estimates are not calibrated enough, hence leading to poor gene perturbations which keep the base model unimproved throughout the cycles.
Finally, the computational time is negligible in our setting, because each sample selection takes less than a minute. That said, GraphReach is around 500 times faster than TypiClust, meaning it scales significantly better with the number of available perturbations. This is important for experiments without predefined candidate perturbations, where the search space can increase from hundreds to tens of thousands.
5 Conclusion
Genomic experiments can not exhaust the available intervention options because they are too costly and time-consuming, thus graph neural networks are utilized to predict the outcomes of these interventions (i.e., gene perturbations). However, the training set for a gene expression model consists of costly genomic experiments itself, calling for more efficient training e.g. based on active learning. Active learning is susceptible to variations during the initial iterations due to the base model’s random initialization. This fact, along with the inherent limitation on the experiment’s horizon creates a risk of wasting genomic experiments in non-informative perturbations.
In this work, we propose graph-based data selection methods that build the train set before the model’s training, such that the risk is alleviated. We develop two methods, GraphReach and MaxSpec, that build the gene perturbation train set using the input knowledge graph. Specifically they select the perturbations that are expected to improve the GNN’s generalization. In the experiments, we see that the data-selection methods perform better then active learning when we have a few hundreds of perturbations, while active learning performs better when we have a few tens. That said GraphReach is the most robust among all the methods, achieving first or second performance in both datasets and both metrics. Taking into consideration these results and that data-selection is devoid of the aforementioned risk, we deem GraphReach a valid approach to train effectively a gene-expression model.
In the future we plan to examine hybrid techniques that utilize graph-based data-selection methods to jump-start the training up to a point and continue with active learning runs for the rest. Moreover, we plan to experiment with multi-gene perturbations and potentially use more involved selection criteria that are common in graph active learning [44, 45].
6 Ethics
This study only involved public datasets that are freely available for academic purposes. There are no obvious ethical considerations regarding negative impacts from the broader application of the method.
{credits}
[…]
The authors have no competing interests to declare that are relevant to the content of this article.
Appendix
Appendix 0.A The GEARS model
GEARS
[34] is a GNN-based model that predicts the gene expressions after a combinatorial perturbation (up to two genes). Let $X\in\mathbb{R}^{N}$ represent the gene expression before perturbation or in the control samples that don’t receive perturbation by definition. Let $P\in[0,1]^{M}$ be the gene perturbation that includes only knockouts (stop the gene=1 or not=0) and finally $Y\in\mathbb{R}^{N}$ be the output gene expression that GEARS predicts. As mentioned in the main text, the cell before perturbation is the same as the control i.e. the control samples have the same gene expression for input and output. Each perturbation includes a number of samples, i.e. $c$ cells that undergo a specific perturbation. We thus have for each perturbation $P$ an input matrix $X\in\mathbb{R}^{c\times N}$ and an output matrix $Y\in\mathbb{R}^{c\times N}$.
The graphs utilized are:
  • Gene coexpression from expressed gene embeddings: An edge is drawn between two genes if they have a pearson correlation in their expressing levels in the training dataset over a predefined threshold.
  • Gene knowledge graph for perturbed gene embeddings: Edges are defined by the Jaccard index in the bipartite gene-pathway from GO (The Gene Ontology), which represents the shared pathways between two genes [9].
It should be noted that gene expression and gene perturbation do not always overlap, meaning we can observe and predict the activity of different genes to the ones we perturb. GEARS learns a function $f(X,P,G_{go},G_{co})=Y$ meaning the post perturbation expressions of all cells that underwent the perturbation $P$. The embeddings are separated between perturbed gene embeddings and expressed gene embeddings. The initial expressed gene embedding is a lookup embedding $f_{g}=\mathcal{Z}\rightarrow\mathcal{R}^{h}$ where $h$ is the hidden dimension size. The GNN embedding of the expressed genes starts with a different lookup embedding and then a simplified convolution [43] over the coexpression graph $G_{co}$.
$x_{g}$ $=ReLU(f_{g}(g))+0.2*ReLU(GNN_{co}(g))$ (6)
$x_{g}$ $=ReLU(W_{g}x_{g})$ (7)
The perturbation index follows a similar lookup $f_{p}=\mathcal{Z}\rightarrow\mathcal{R}^{h}$, and a pass to the respective $GNN_{p}$ similar to the aforementioned model. If it is a control sample then this is overlooked. The perturbation embeddings for each perturbation in the sample, represented by an array $[x_{p}^{1},..,x_{p}^{l}]$, where $l$ is the number of genes perturbed, are fused through an MLP:
$x_{p}=ReLU(W_{fuse}\sum_{i}^{l}x_{p}^{i})$ (8)
The perturbation embedding is then added to each of the expressed gene embeddings $j$:
$x_{fused}^{j}=x_{g}^{j}+x_{p}$ (9)
followed by another MLP, that gives us the final post perturbation gene embedding that contains information from both graphs:-
$x_{post}^{j}=ReLU(W_{fused}x_{fused}^{j})$ (10)
Before adding this to the current gene expression as observed in the cell, the model renders the post-perturbation embedding of each expressed gene into a scalar $z_{comb}^{j}\in\mathcal{R}$ :
$z_{comb}^{j}=w_{comb}^{j}x_{post}^{j}+b_{comb}^{j}$ (11)
Subsequently, they form $z=[z_{comb}^{1},...,z_{comb}^{N}]\in\mathcal{R}^{N}$ for all genes in the cell which is passed through an MLP.
$h_{cg}=ReLU(w_{cg}z)$ (12)
The idea is to use the ’cross gene embedding’ $h_{cg}$ to capture the state of all expressed genes when predicting the individual gene expression since genes interact with each other. It is concatenated with the original $z_{comb}^{j}$ to derive the final perturbation effect for each gene.
$\hat{z}_{j}=w_{cg}[z_{comb}^{j}||h_{cg}]+b_{u}$ (13)
Then $\hat{z}$ is added to the original gene expression to get the prediction.
$Y=X+\hat{z}$ (15)
The main cost function is:
$L=\frac{1}{T}\sum^{T}_{k=1}\frac{1}{T_{k}}\sum^{T_{k}}_{l=1}\frac{1}{K}\sum^{K}_{u=1}(Y_{u}-\hat{Y}_{u})^{(2+\gamma)}$ (16)
where $T$ is the batch of perturbations, $k$ is a perturbation and $T_{k}$ are the corresponding cells (samples) and each cell has $K$ gene expressions. The cost function that calculates the direction is
$L_{d}=\frac{1}{T}\sum^{T}_{k=1}\frac{1}{T_{k}}\sum^{T_{k}}_{l=1}\frac{1}{G}\sum^{K}_{u=1}[sign(y_{u}-y^{c}_{u})-sign(\hat{y}_{u}-y^{c}_{u})]^{2}$ (17)
where $G$ is the number of perturbations, $y^{c}_{u}$ is the average gene expression for all control samples. The control samples have the same $y$ and $x$, so $y^{c}_{u}=x^{c}_{u}$ but the pre-perturbational values $x_{u}$ which can be considered also control are not utilized. The final cost is $L+\lambda L_{d}$.
Note that during training, all genes with zero expression are not taken into account for the calculation of loss. Uncertainty is computed through an evidential layer (predicts the expression and a log variance $\sigma$ for each gene). The main evaluation (in train and test) is the MSE between the predicted postperturbation expression and the real one, for the top 20 most deferentially expressed genes. The pearson correlation between the whole vectors of gene expression is also computed.
Appendix 0.B Proof of Submodularity
Following the definitions of the main paper we know that for the set function of reachability $R$
  • $R(A\cup\{u\})=R(A)\cup R(\{u\})$
  • $R(B)=R(A)\cup(R(B)\setminus R(A))$
Then we can start from the definition of submodularity:
$R(A\cup\{u\})-R(A)\supseteq$ $R(B\cup\{u\})-R(B)\longleftrightarrow$
$R(A)\cup R(\{u\})-R(A)\cup(R(B)\setminus R(A))\cup R(\{u\})$ $+R(B)-R(A)\supseteq\emptyset\longleftrightarrow$
$R(A)\cup R(\{u\})-(R(A)\cup R(\{u\}))\cup(R(B)\setminus R(A))$ $+R(B)-R(A)\supseteq\emptyset$
Where we can break the proof into cases:
if $(R(B)\setminus R(A))\cap R(\{u\})=\emptyset$ or $(R(B)\setminus R(A))\supseteq R(\{u\})$ then we have
[…]
$R(B)\supseteq R(A)+R(B)$ $\setminus R(A)$
which is the definition above.
if $(R(B)\setminus R(A))\subseteq R(\{u\})$ then we have
[…]
which is correct from monotonicity $R(B)\supseteq R(A)$
Finally iff $(R(B)\setminus R(A))\cap R(\{u\})\neq\emptyset$ then we have
[…]
$R(B)+R(\{u\})\supseteq R(A)+R(\{u\})\cup(R(B)\setminus R(A))$
which is an equality.
Hence submodularity stands $|R(A\cup\{u\})-R(A)|\geq|R(B\cup\{u\})-R(B)|$.

### v3

@@
 While typical active learning mitigates this issue over many iterations, the limited number of experimental cycles in genomic studies exacerbates the risk.
-To this end, we propose graph-based one-shot data selection methods for training gene expression models.
-Unlike active learning, one-shot data selection predefines the gene perturbations before training, hence removing the initialization bias.
-The data selection is motivated by theoretical studies of graph neural network generalization.
+To this end, we propose graph-based data filtering as an alternative.
+Unlike active learning, data filtering selects the gene perturbations before training, meaning it is free of bias due to random initialization and initial random selection.
+Moreover, reducing the iterations between the wet lab and the model provides several operational advantages resulting in significant acceleration.
+The proposed methods are motivated by theoretical studies of graph neural network generalization.
 The criteria are defined over the input graph and are optimized with submodular maximization.
-We compare them empirically to baselines and active learning methods that are state-of-the-art on this problem.
-The results demonstrate that graph-based one-shot data selection achieves comparable accuracy while alleviating the aforementioned risks.
+We compare them empirically to baselines and active learning methods that are state-of-the-art.
+The results demonstrate that graph-based data filtering achieves comparable accuracy while alleviating the aforementioned risks.
 Keywords:
-Active Learning Graph Neural Networks Gene Perturbation Prediction
+Gene Perturbation Prediction Active Learning Optimal Experimental Design
 1 Introduction
@@
 Techniques such as CRISPR interference (CRISPRi) and PerturbSeq [5] have revolutionized the landscape of genomic experimentation by enabling high-throughput screenings.
-However, analyzing combinatorial gene perturbations remains challenging [38, 28], leading most studies to focus on single-gene perturbations.
-Moreover, most CRISPR-based PerturbSeq experiments are restricted to the order of a few hundreds of gene perturbations [10, 1, 31] due to budget and time constraints, despite the availability of over 20,000 potential gene targets.
-To complement experimental observations, several machine learning models have been developed to predict the outcome of gene perturbations on a given cell  [34, 7, 8, 13, 22, 23, 14].
+However, analyzing combinatorial gene perturbations remains challenging [39], leading most studies to focus on single-gene perturbations.
+Moreover, most CRISPR-based PerturbSeq experiments are restricted to the order of a few hundreds of gene perturbations [10, 1, 33] due to budget and time constraints, despite having over 20,000 potential gene targets.
+To assist exploring this vast space, machine learning models have been developed to predict the outcome of gene perturbations on a given cell  [36, 7, 8, 14, 25, 26, 15].
+These models leverage diverse neural architectures, ranging from autoencoders [7] to graph neural networks [36], which require a substantial number of training samples to achieve generalization.
+These labeled samples stem from genomic experiments, which, as mentioned above, are rather costly.
+This, coupled with the fact that the models are disease specific, meaning they are trained on single cell data from the same hypotheses, render the model training an optimization problem in and of itself.
+Typically this is addressed with active learning [38] which defines an iterative interaction between the wet lab and the model in order to retrieve the genes that will optimize the training, as shown in Fig.
+1(a).
+Active learning strategies can be based on i) uncertainty, which query instances where the model is least confident, ii) information theory, which seek instances that are expected to provide the most information about the model parameters and iii) density, which ensures that selected samples are representative of the input distribution.
+The first two strategies use the model trained on the available data to guide the selection at each iteration.
+On the contrary, density-based methods typically rely on the dataset distribution without input from the model, which, however, is not possible in the current setting.
+Specifically, in SOTA models such as GEARS [36], the only input available before executing the genomic experiment is the candidate gene perturbation name and a knowledge graph defined a priori.
+Hence, density-based methods rely on the model to map the genes to a learnable embedding space and then act upon it [18].
+These renders all our available strategies model-based, which is highly problematic.
 (a) Active learning.
-(b) One-shot data selection.
-Figure 1: The difference between active learning and one-shot data-selection for genomic experiments with GEARS [34].
-These models leverage diverse neural architectures, ranging from autoencoders [7] to graph neural networks [34], and they require a large number of training samples to generalize effectively.
-This reliance on supervision occasionally defeats their main purpose (i.e., minimizing the experimental cost), because the labels are themselves retrieved from genomic experiments.
-Additionally, disease-specific variations in gene expression make it difficult to train a universal model, necessitating the development of models tailored to specific experimental conditions.
-Active learning [37] offers a promising solution by enabling iterative interaction between the wet lab and the machine learning models, to optimize data collection and hence training, with a limited number of cycles, as shown in Fig.
-1(a).
-In the context of CRISPRi-based PerturbSeq experiments, these cycles tend to be restricted (e.g.
-5 cycles), as each experiment in the lab may take up to a month [17].
-Given the vast number of possible genes to explore, active learning strategies are well-suited to efficiently train predictive models while minimizing the experimental costs.
-However, active learning depends on the model’s initialization, which poses a non-negligible risk in this context.
-Fig.
-2 illustrates the amount of variation induced in the genomic perturbations selected to train the GEARS model [34], based on 5 different random initializations.
+(b) data filtering.
+Figure 1: The difference between active learning and data filtering for genomic experiments with GEARS [36].
+[…]
+To quantify the sensitivity of the framework to initialization bias, Fig.
+1 illustrates the amount of variation induced in the gene perturbations selected to train GEARS, based on 5 different random initializations of the model using the SOTA algorithm for the problem [18].
 In particular, it illustrates the number of genes that are selected in any one given cycle as black bars and the total number of unique genes that are selected for the 5 differently initialized models as the orange bars.
-Since the orange bars are often more than twice as large than the black ones, we observe that the perturbed genes differ significantly for each initialization, i.e., an extreme sensitivity of the selected genes to the model’s initialization.
-The effect is more pronounced in the initial cycles, as the models become more stable in subsequent cycles and converge to similar choices.
-However, given the aforementioned limitation in the cycle horizon, the train set may mistakenly include gene perturbations that are difficult to correct afterwards.
-In typical active learning scenarios, a longer horizon leads to an overall stable selection, mitigating early-stage errors over time.
-Nevertheless, in this case, the wet lab may risk losing months of work and resources for a model that is not properly initialized and may not converge in time.
+Since the orange bars are often more than twice as large as the black ones, we observe that the perturbed genes differ significantly for each initialization, i.e., an extreme sensitivity of the selected genes to the model’s initialization.
+The effect is more pronounced in the initial cycles because the models become more stable in subsequent cycles and converge to similar choices, as expected by an AL procedure.
+However, given the aforementioned limitation in the cycle horizon, the train set does not recover.
 Figure 2: The total number of unique genes perturbed in active learning experiments in two datasets with 5 different random seed initializations of GEARS [36], on the same train-test split.
-Compared to the budget for gene perturbations, i.e., the number of genes to be perturbed in each cycle, this indicates the variation associated with building the dataset based on the model in the initial cycles.
-To mitigate this risk, we explore the use of one-shot data selection methods as an alternative to active learning for the problem of training gene expression predictors.
-By selecting the samples upfront based on the input of the model, as seen in Fig.
-1(b), we can provide competitive results while exhibiting:
+Compared to the budget for gene perturbations, i.e., the number of genes to be perturbed in each cycle, this indicates the sensitivity of the gene selection process to the model initialization.
+To mitigate this risk, we explore methodologies that disentangle the model with the selection strategy.
+Specifically, we focus on data filtering methods as an alternative to active learning [22].
+Data filtering selects samples before training the model, based on the information contained a priori.
+By selecting the samples upfront, as seen in Fig.
+1(b), we can provide competitive results while exhibiting several key advantages:
   • Independence of model initialization.
-A data-selection method runs before the model’s training and hence it does not depend on the model’s predictions.
-This reduces the risk of wasting experimental resources on erroneous gene perturbations.
+Data-filtering does not depend on the model’s predictions.
+This reduces the risk of wasting experimental resources on erroneous gene perturbations due to lack of training data, random model initialization and reduced iteration number.
+  • Decrese experimental time.
+In active learning, each iteration requires the wet lab to run experiments, which may last weeks each.
+In contrast, with data-filtering, we ask for the same number of experiments but only once.
+Since the cell assays can run in parallel, this means data filtering will be solved roughly in the time required for one iteration of active learning.
+This translates to 3-4 months of acceleration in real life.
   • Reduced operational delay.
-Each iteration in active learning introduces extra computational overhead: model retraining, recalculating uncertainty scores, etc.
-Moreover iterative experiments exhibit delays due to logistical complexities and communication, compared to one large experiment.
-These problems can add a substantial delay, which is not present in one uniform experiment.
-We propose two one-shot data selection methods for training the state-of-the-art GEARS model [34], although our methods can work with any graph-based predictor.
+Each iteration requires a computational overhead for model retraining, computing the acquisition step, etc.
+Moreover, iterative experiments exhibit non-negligible delays due to logistical complexities and communication between ML team and the wetlab.
+These problems can add a substantial delay, which is not present when we only run one experiment.
+We propose two data filtering methods for training the GEARS model, although our methods can work with any graph-based predictor.
 We rely on theorems for GNN generalization to define proxy criteria to guide the sample selection based on GEARS’ input knowledge graph.
-The criteria are optimized greedily to build the train set.
+The criteria are optimized greedily to build the training set.
+Note that, we focus solely on the knowledge graph because the coexpression network is derived from the cell assays, which by default are not available before the experiment takes place.
 Our contributions are as follows:
-  • Two novel data-selection criteria based on generalization theorems for GNNs that utilize the input knowledge graph and are model-free.
-  • A cost effective lazy forward submodular maximization algorithm to optimize one of the criteria, and a respective greedy heuristic for the other.
-  • An empirical comparison including two active learning methods (one of which is SOTA in our problem) and two baselines, indicating the effectiveness of one-shot data selection.
+  • Two novel data-filtering criteria based on generalization theorems for GNNs that utilize the input knowledge graph and are model-free.
+  • A submodular maximization algorithm to optimize one of the criteria, and a respective greedy heuristic for the other.
+  • An empirical comparison including two active learning methods (one of which is SOTA in our problem) and two baselines, indicating the effectiveness of data filtering.
 2 Related Work
@@
 However, prior knowledge is not guaranteed to exist for every dataset.
-Our one-shot data-selection methodology addresses this gap along with the fundamental problems of the active learning analyzed in sec.
+Our data filtering methodology addresses this gap along with the fundamental problems of the active learning analyzed in sec.
 1.
 We focus on training GEARS effectively while lacking prior information and the ability to take feedback from the model.
-The former is a plausible assumption since data for specific hypothesis are not that commonly available, and the latter is motivated by the variation inherent in the initial cycles of an active learning process and the cost to acquire samples in our context.
+The former is a plausible assumption since data for specific hypotheses are not that commonly available, and the latter is motivated by the variation inherent in the initial cycles of an active learning process and the cost to acquire samples in our context.
 3 Methodology
-In this section we first formulate the problem and clarify the motivation for pure graph-based methods for one-shot data selection.
-Subsequently we introduce the two methods for one-shot data selection.
+In this section, we first formulate the problem and clarify the motivation for graph-based approaches, and subsequently we introduce the two methods for data filtering.
 3.1 Formulation
-We are given a gene graph $G=(V,E),$ with $|V|=N$ and $|E|=M$, where some of the nodes are candidate perturbations.
-We assume that the dataset is partially labeled, meaning there is a subset of nodes $\mathcal{L}\subset V$ labeled.
-Each label is a matrix $Y_{u}\in\mathcal{R}^{c\times|g|}$ which represents the gene expression of a set of genes $g$ (which can belong to $V$ or not) of $c$ cells that underwent perturbation $u\in V$, along with the respective pre-perturbation gene expressions $X_{u}\in\mathcal{R}^{c\times g}$.
-There is respectively a set of genes $\𝑈\𝑉U=V\backslash\mathcal{L}$ that are unlabeled.
-GEARS learns how to predict $Y_{u}$ based on a given perturbation $u$.
-The problem we focus on is:
-    Iteratively build the train set $S$ with the $b$ best genes to perturb from set $U$ to maximize GEARS’ accuracy on an unseen set of gene perturbations.
-We aim to build the train set $S$ in order to cover as much and as diverse information as possible to train the model.
-Similar to the core-sets [36], we define a covering function that maximizes the model’s reach, but instead of defining it in the embedding space, we define it in the knowledge graph.
-The main reason behind this is that our graphs are not accompanied by node features.
-The node embeddings are randomly initialized and learned through training.
-This means any density-based method would be conditioned on the model initialization problem we explained in sec.
-1.
-To be specific, in GEARS and most gene expression models, the pre-perturbation gene expressions $X_{u}$ are not utilized during the prediction process as node features, but rather they are added to the predicted perturbation to arrive at $Y_{u}$.
-This happens because the prediction of the perturbation should be independent to the state of the cell $X_{u}$, as in real experiments, the cells will not exist until we have already decided which perturbation to run.
-This is why gene expressions models predict directly the gene expression displacement given the name of the perturbation, without using the cell $X_{u}$ [34, 7, 13].
-Hence, in the absence of input node features and with the intention to remove the dependence of the train set selection on the model (due to the aforementioned variability in the resulting election), we are left with the knowledge graph as our sole remaining source of information to select the train set.
-The GEARS architecture is described in detail in the appendix for reference and further clarification.
-3.2 GraphReach
-To understand how to use the graph to guide our selection, we start with the graph convolution layer of GEARS. The Simplified Graph Convolution (SGC) layer [43] is defined as
-$\tilde{\mathbf{A}}=\mathbf{D}^{-1/2}\mathbf{A}\mathbf{D}^{-1/2}$ (1)
-$\mathbf{H}_{t+1}=\tilde{\mathbf{A}}^{K}\mathbf{H}_{t}\mathbf{W}$ (2)
-where $\mathbf{H}_{t}$ are the node representations at layer $t$, $\mathbf{W}$ represents the learnable parameters of the model, $\mathbf{A}$ is the graph’s adjacency matrix, $K$ is the power of the adjacency predefined as 1 in GEARS, and $\mathbf{D}$ is the diagonal degree matrix.
-It should be noted that SGC is known to exhibit better generalization gaps than most of the other layers [18, 41], rendering it a well-suited module for our task.
-Intuitively, our main goal is to label nodes such that we cover as much information as possible from the input, akin to core-sets.
-As mentioned above, the SGC aggregates information from the whole network; yet only the representations of nodes that are within the reach of labeled training nodes will get updated through the message passing layers.
+We are given a gene knowledge graph $G=(V,E)$, where $|V|=N$ and $|E|=M$, and each node represents a gene, some of which correspond to candidate gene perturbations.
+Our goal is to select a subset $S\subset V$, under a budget constraint $|S|=b$ with $b\ll N$, to experimentally perturb in a wet lab setting.
+For each selected gene $u\in S$, the perturbation experiment yields paired measurements of pre- and post-perturbation gene expression: $X\in\mathbb{R}^{c\times g}$ for $c$ single-cell profiles and $g$ expressed genes (not to be confused with the perturbed genes), and corresponding post-perturbation outputs $Y\in\mathbb{R}^{c\times g}$.
+Thus, the perturbation effect is modeled as an additive change $\hat{P}_{u}\in\mathbb{R}^{g}$ such that $\hat{Y}_{u}=X+\hat{P}_{u}$, and the model learns a mapping $M:V\rightarrow\mathbb{R}^{g}$.
+Note that the perturbation is run on multiple cells with varying gene expressions.
+The model $\theta^{S}$, trained only on the selected set $S$, is a graph neural network (GNN) such as GEARS. The loss function is the mean squared error between predicted and observed expression:
+$\ell(X,Y,u;\theta^{S})=\|Y-(X+\hat{P}_{u})\|^{2},$ (1)
+where $\hat{P}_{u}=M(u;\theta^{S})$.
+The problem we address is to select the most informative subset of genes to perturb, such that the trained model generalizes well to unseen perturbations $U\subseteq V\setminus S$.
+Formally, we aim to:
+$\underset{S\subset V,\ |S|=b}{\arg\min}\ \sum_{u\in U}\ell(X_{u},Y_{u},u;\theta^{S})\leavevmode\nobreak\ ,\leavevmode\nobreak\ s.t.|S|=b,$ (2)
+where $X_{u},Y_{u}$ denote the inputs and targets for perturbation $u$, and $\theta^{S}$ is trained only on data from $S$.
+3.2 Motivation
+As mentioned above, in normal density-based active learning we try to cover as much of the input space as possible.
+For example, in core-sets[37], it is proven for i.i.d.
+data that if the training samples are selected such that each test sample is less then $\delta$ away from one of them, we are guaranteed to reduce generalization error with a factor analogous to $\delta$ and the total number of samples.
+This can be seen as ’covering’ the covariate space with the training samples that have a $\delta$-radius.
+Their main assumption is that samples that are close in space exhibit similar outcomes.
+In our case this assumption translates to homophily ,i.e.
+, the effects of genes that are close in the KG are similar, and the $\delta$ radius can be defined based on hop distance.
+However, since GNNs utilize message passing, they break the i.i.d.
+assumptions which makes this result unusable.
+In terms of active learning for GNNs, labeling a node incurs an information gain to every other node in their receptive field [43, 46, 45] (,i.e, $k$-hop neighborhood), meaning that the supervision signal propagates due to message passing.
+In the same vain, [19] argue that training is improved by maximizing the diversity of aggregated features, which translates to maximizing the number of unique nodes in the training.
+The problem with these approaches is that the metrics depend on the node distances, the features of unlabeled nodes, and the representations of the model.
+In our setting, $X$ and $Y$ are available only for labeled nodes, and we have no access to the model for the reasons detailed in section 1, hence we can not directly utilize this framework.
+These approaches agree with the basic intuition that although a GNN aggregates information from the whole network, the only representations that will get updated correspond to the nodes that are within the reach of the training nodes.
 For example, if there are $L$ layers, message passing reaches up to the $L$-hop neighborhood of labeled nodes in set $S$.
@@
 Thus, to maximize the reach of the supervision signal and cover as much of the input information as possible, we should aim to maximize the number of nodes reached by the training set $S$.
-Beyond our intuition, this is supported by theoretical insights into the generalization capability of GNNs.
-The number of nodes that we have to predict and the number of samples that we can add in each cycle is fixed, so we can not change the VC dimension [35].
-However, theoretical studies suggest that the generalization of a GNN is improved based on the distance between the training and the test nodes in the graph [25, 46], supporting our intuition that covering a larger part of the graph with training nodes-and thereby reducing the distance to the test set- is beneficial to the model.
-Note that these works define the distance as the structural and distributional difference; nevertheless, since we lack node features, the structural distance is the sole factor in our case.
-Based on the above, we can define our criterion function, similar to an acquisition function in active learning, such that we maximize the number of nodes reached by the set $S$.
+Beyond the aforementioned literature, this is supported by theoretical insights into the generalization capability of GNNs.
+Theoretical studies suggest that the generalization is improved as we reduce the distance between the training and the test nodes in the graph [28, 48].
+This supports our intuition that covering a larger part of the graph with training nodes-and thereby reducing the distance to the test set- is beneficial to the model.
+Note that these works define the distance as the structural and distributional difference; nevertheless, since we do not have access to the features or the model, we will define this covering functions on the only input space we have, the knowledge graph.
+Figure 3: An overview of the data filtering methodology.
+Our sole input is the knowledge graph from GO[36] and a list of candidate perturbations.
+The data filtering algorithm ,e.g.
+, GraphReach selects the set of gene perturbations and they are given to the wet lab for the experimental part.
+Finally, the single cell gene expression data is given to GEARS for training and validation.
+3.3 GraphReach
+We define our criterion function, similar to an acquisition function in active learning, such that we maximize the number of nodes reached by the set $S$.
 Specifically for a new node $v$, the criterion is defined as the additional reachability that $v$ adds in the set of train perturbations $S$
@@
 where $R_{L}$ is the set of nodes reachable from set $S$, based on the number of SGC layers $L$ in the definition of GEARS
-$R_{L}(S,A)=\{v\in V\mid(\mathbf{A}^{\ell}_{uv}>0\text{ for some }u\in S,\text{and }1\leq\ell\leq L\}$ (4)
-The subscript $L$ is omitted in subsequent discussion as it remains stable throughout all experiments in GEARS.
+$R_{L}(S,A)=\{v\in V\mid\mathbf{A}^{\ell}_{uv}>0\text{ for some }u\in S,\text{ and }1\leq\ell\leq L\}$ (4)
+The subscript $L$ is omitted in subsequent discussion as it remains stable throughout all experiments.
 The criterion function selects greedily the node $v$ maximizing $\alpha(v,S)$ in every iteration.
@@
 We employ a cost effective lazy forward approach to accelerate the acquisition step without loss of accuracy [24].
-The final data selection algorithm, called GraphReach from graph reachability, is shown in Alg.
-1.
+The final data filtering algorithm, called GraphReach from graph reachability, is shown in Alg.
+1, and an overview of the overall step-by-step data filtering approach can be seen in Fig.
+3.
 Algorithm 1 GraphReach
@@
 The final graph contains over 9,800 nodes and 200,000 edges.
-It should be noted that GEARS utilizes the whole graph to learn representations and uses the gene perturbation datasets for supervision only.
+It should be noted that GEARS utilizes the whole knowledge graph to learn representations and uses the gene perturbation datasets for supervision only.
 In other words, the model uses the perturbations corresponding to the train set to compute the loss, but the SGC layer runs over the whole graph, including the test perturbations and unrelated genes, in a semi-supervised learning fashion [21].
-The data selection methods follow suit and utilize the whole graph.
-For example the reachability of a gene perturbation in Replogle is not defined only on the 1081 perturbations, but on the 9,800 nodes.
+The data filtering methods follow suit and utilize the whole graph.
+For example, the reachability of a gene perturbation in Replogle is not defined only on the 1088 candidate perturbations, but on the whole 9,800 nodes.
 Similar to  [18], we start with a random test set that contains roughly 10% of the available perturbations, i.e., 100 for Replogle and 8 for Adamson.
@@
 As can be seen in Tab.
-1, the models developed from the data selection methods perform better than SOTA in one dataset and close to it in another.
+1, the models developed from the data filtering methods perform better than SOTA in one dataset and close to it in another.
 To be precise, while MaxSpec and TypiClust are the top methods in Reglogle and Adamson respectively, GraphReach has overall the most balanced performance being first or second in all datasets and metrics with minimal differences (considering the standard deviations) from the top.
 Fig.
-3(a) and 4(a) show that GraphReach outperforms marginally TypiClust in 3 out of 5 cycles in the Adamson dataset.
+4(a) and 5(a) show that GraphReach outperforms marginally TypiClust in 3 out of 5 cycles in the Adamson dataset.
 Both achieve a better accuracy by the third cycle than the other models do in 5 cycles, with the exception of ACS-FW at the 5th cycle.
@@
 TypiClust $155\pm 6$ $\mathbf{138\pm 47}$ $919\pm 7$ $\mathbf{963\pm 9}$
-Data Selection MaxSpec $\mathbf{150\pm 4}$ $160\pm 49$ $920\pm 9$ $958\pm 11$
+Data Filtering MaxSpec $\mathbf{150\pm 4}$ $160\pm 49$ $920\pm 9$ $958\pm 11$
 GraphReach $151\pm 4$ $143\pm 44$ $\mathbf{921\pm 8}$ $962\pm 9$
@@
 Best method denoted with bold and second best is underlined.
-For the Replogle dataset, the active learning methods are outperformed clearly by the graph-based data selection methods and even by random selection in the initial cycles.
+For the Replogle dataset, the active learning methods are outperformed clearly by the graph-based data filtering methods and even by random selection in the initial cycles.
 ACS-FW in particular has considerably worse performance than random, which is why it is not included in Fig.
-3(b) and Fig.
-4(b), while TypiClust consistently falls behind the remaining methods.
+4(b) and Fig.
+5(b), while TypiClust consistently falls behind the remaining methods.
 In contrast, MaxSpec is the best method in Replogle but argueably the worst in Adamson.
@@
 all available genes), we expect this effect to be greater.
-This further supports the need for the data selection methods, which have no cycle of deteriorating accuracy.
+This further supports the need for the data filtering methods, which have no cycle of deteriorating accuracy.
 Additionally, the results are aligned with the literature that confirms random selection is stronger than most active learning in limited cycles [40, 4, 30, 49].
@@
 This fact, along with the inherent limitation on the experiment’s horizon creates a risk of wasting genomic experiments in non-informative perturbations.
-In this work, we propose graph-based data selection methods that build the train set before the model’s training, such that the risk is alleviated.
+In this work, we propose graph-based data filtering methods that build the train set before the model’s training, such that the risk is alleviated.
 We develop two methods, GraphReach and MaxSpec, that build the gene perturbation train set using the input knowledge graph.
 Specifically they select the perturbations that are expected to improve the GNN’s generalization.
-In the experiments, we see that the data-selection methods perform better then active learning when we have a few hundreds of perturbations, while active learning performs better when we have a few tens.
+In the experiments, we see that the data filtering methods perform better then active learning when we have a few hundreds of perturbations, while active learning performs better when we have a few tens.
 That said GraphReach is the most robust among all the methods, achieving first or second performance in both datasets and both metrics.
-Taking into consideration these results and that data-selection is devoid of the aforementioned risk, we deem GraphReach a valid approach to train effectively a gene-expression model.
-In the future we plan to examine hybrid techniques that utilize graph-based data-selection methods to jump-start the training up to a point and continue with active learning runs for the rest.
+Taking into consideration these results and that data filtering is devoid of the aforementioned risk, we deem GraphReach a valid approach to train effectively a gene-expression model.
+In the future we plan to examine hybrid techniques that utilize graph-based data filtering methods to jump-start the training up to a point and continue with active learning runs for the rest.
 Moreover, we plan to experiment with multi-gene perturbations and potentially use more involved selection criteria that are common in graph active learning [45, 47].
@@
 The main cost function is:
-$L=\frac{1}{T}\sum^{T}_{k=1}\frac{1}{T_{k}}\sum^{T_{k}}_{l=1}\frac{1}{K}\sum^{K}_{u=1}(Y_{u}-\hat{Y}_{u})^{(2+\gamma)}$ (16)
+$L=\frac{1}{T}\sum^{T}_{k=1}\frac{1}{T_{k}}\sum^{T_{k}}_{l=1}\frac{1}{K}\sum^{K}_{u=1}(Y_{u}-\hat{Y}_{u})^{2}$ (16)
 where $T$ is the batch of perturbations, $k$ is a perturbation and $T_{k}$ are the corresponding cells (samples) and each cell has $K$ gene expressions.